Vertex (7,11), directrix x-1 standard equation of parabola
1
Expert's answer
2021-10-11T05:48:13-0400

The equation of the parabola has the form "(y-b)^2=2p(x-a)" , where "(a;b)" is a vertex of parabola, so "a=7, b=11". "x=-\frac{p}{2}" is a directrix . Vertex is halfway between focus and direcrix. The focus lies to the right side of the directrix, so the parabola opens to the right. As "x=1" , so "p=7-1=6". The focus of the parabola has coordinates "(12;0)". So the standard equation of parabola is "(y-11)^2=12(x-7)"